OpenStudy (anonymous):

The table shows the results of drawing one marble from a bag of colored marbles. Each marble was replaced after the outcome was recorded. If there are a total of 50 marbles in the bag, about how many of the marbles are green? Color Number of Pulls red 6 green 4 orange 7 black 3 A. 10 B. 4 C. 8 D. 20

OpenStudy (anonymous):

pls help me

OpenStudy (anonymous):

i ill fan and medal

OpenStudy (anonymous):

im not really sure but my guess is 10 green marbles. not 20 b/c the amount of marbles picked is less and not 4 b/c thats how many were picked out of 50 so either 8 or 10 and 10 seems reasonable b/c its out of 50.

OpenStudy (anonymous):

is the answer 10

OpenStudy (ranga):

Yes.

ganeshie8 (ganeshie8):

as the number of trials increase, the proportion comes closer to probability. so setup a proportion : \(\large \frac{4}{20} = \frac{x}{50}\)

ganeshie8 (ganeshie8):

solve \(x\)

ganeshie8 (ganeshie8):

10 is \(\large \color{red}{\checkmark}\)

OpenStudy (anonymous):

yup. i like that

OpenStudy (anonymous):

P(green)= 4/20 = 1/5 total no. of green marbles= (1/5) x 50 = 10